# Knowledge Independence Breeds Disruption but Limits Recognition

### v1

@@
Abstract
Recombinant growth theory highlights the pivotal role of cumulative knowledge in driving innovation  (?, ?, ?, ?). Although interconnected knowledge facilitates smoother dissemination  (?, ?), its connection to scientific disruption remains poorly understood. Here, we quantify knowledge dependence based on the degree to which references within a given paper’s bibliography cite one another. Analyzing 53.8 million papers spanning six decades, we observe that papers built on independent knowledge have decreased over time. However, propensity score matching and regression analyses reveal that such papers are associated with greater scientific disruption, as those who cite them are less likely to cite their references. Moreover, a team’s preference for independent knowledge amplifies its disruptive potential, regardless of team size, geographic distance, or collaboration freshness. Despite the disruptive nature, papers built on independent knowledge receive fewer citations and delayed recognition. Taken together, these findings fill a critical gap in our fundamental understanding of scientific innovation, revealing a universal law in peer recognition: Knowledge independence breeds disruption at the cost of impact.
With the shift in modern science toward large-scale, complex endeavors  (?, ?, ?), and the emergence of new scientific paradigms  (?), there is an increasing need to integrate diverse knowledge streams in scientific research  (?, ?). Recombinant growth theory  (?) posits that a larger body of knowledge provides more opportunities for novel combinations, thereby fostering greater innovation. However, empirical observations seem to contradict this prediction, as innovation has paradoxically slowed down even declined over time  (?, ?). This suggests that simply combining existing knowledge is insufficient to drive innovation, and that the way such knowledge is combined matters  (?).
Previous studies have analyzed the research strategies underpinning scientific papers, highlighting a widespread tendency to rely on well-established knowledge  (?, ?). This tendency extends to individual career trajectories, where scholars tend to exploit familiar topics rather than explore unfamiliar ones  (?, ?). Yet, innovative insights seem to emerge from unfamiliar topics, particularly through papers that draw on atypical combinations of knowledge  (?). Prior research has shown that novel combinations of research contents  (?), datasets  (?), methods  (?, ?), and references  (?, ?) are associated with greater scientific impact. Beyond conventional impact metrics, recent studies have linked such knowledge recombination to scientific disruption, i.e., the degree to which a paper overshadows the prior work it builds upon  (?, ?). Specifically, disruption has been associated with the novel combinations of references  (?) and research methods  (?).
Despite these numerous studies, one critical aspect of knowledge combination remains largely unexplored in the Science of Science literature to date—the notion of path dependence  (?). Intuitively, path dependence captures the idea that current and future decisions, outcomes, or developments in a system are shaped by the “path” (or sequence) of events or choices that preceded them  (?, ?, ?). In technical domains, this notion has been extensively studied, particularly in the context of patents. Research has shown that path dependence, the reliance on densely-interrelated patents when creating new ones, may foster incremental contributions by building on established technologies, reducing development costs, and promoting efficient progress  (?, ?, ?). However, it may also constrain innovation by locking inventors into existing trajectories, making it more challenging to explore novel or disruptive ideas outside established frameworks  (?, ?, ?). What remains unclear is whether this general law—that path dependence consolidates cumulative knowledge while path independence encourages innovations—extends from technology (i.e., patents) to science (i.e., papers).
Addressing the aforementioned gap would require a measure of path dependence that can be applied to the context of Science of Science. However, existing measures are predominantly qualitative, offering a dichotomous view in which paths are either dependent or independent, rather than a continuous perspective encompassing a spectrum of path dependence to independence  (?, ?, ?). The only continuous measures proposed in the literature operate at the meso-level, assessing the extent to which technical institutions rely on their past knowledge and experience  (?, ?). Yet, developing a continuous measure of path dependence or independence at the micro-level remains an open question.
Here, we introduce a formal definition of knowledge path independence in the Science of Science. Intuitively, this measure quantifies the extent to which the references in a paper are “independent” in the sense that they do not cite one another. For instance, if reference $A$ cites reference $B$, then this would be a knowledge path from $B$ to $A$. As such, knowledge path independence decreases as these paths grow longer and more frequent, signifying that the paper is built on a well-knitted body of knowledge. Throughout the remainder of this article, we will omit the term “path” and simply write “knowledge independence” to avoid potential confusion with the broader notion of paths in graph theory.
Armed with our measure of knowledge independence, we analyze 53.8 million publications from 1955 to 2017 in the Web of Science, and uncover a marked drop in knowledge independence across disciplines. At the same time, knowledge independence is becoming increasingly associated with scientific disruption. This relationship remains robust even when controlling for focal papers’ impact and novelty, their references’ age, impact, and disruption, as well as the composition of the team that creates the paper. Propensity score matching indicates a significant effect of knowledge independence on scientific disruption, while normalized regression analyses indicate that this effect is greater than those of the above factors. We further uncover the mediating role of knowledge independence in the relationship between disruption and team composition, regardless of the team size, geographic distance, or collaboration freshness. Finally, we find that papers built on independent knowledge receive fewer citations and delayed recognition, thereby explaining the negative association between disruption and impact. We confirm the robustness of all these findings when examining a different dataset focusing on publications, namely, SciSciNet  (?). Overall, our findings address a crucial gap in the fundamental understanding of scientific innovation, demonstrating that the recombination of independent knowledge breeds disruption, thereby becoming a knowledge broker that occupies a niche in the intellectual ecosystem. Yet, this very independence simultaneously increases the challenge of knowledge diffusion, leading to limited and delayed recognition.
@@ Abstract
To quantify knowledge independence in the context of bibliometrics, we develop a measure that relies on the citation network. This is because citations serve as tangible indicators of knowledge flow, reflecting the dependency of the citing paper on its references. Intuitively, for any given paper, the measure quantifies the degree to which its references are “independent” in the sense that they do not cite one another. To this end, let us introduce two types of references: An $i$-type reference is one that does not cite any other work within the same reference list, whereas a $j$-type reference is one that does cite at least one other work in this list. Then, knowledge independence (KI) is measured as the difference between the fraction of $i$-type and $j$-type references. More formally:
$\text{KI}=\frac{n_{i}-n_{j}}{n_{i}+n_{j}},$ (1)
@@
where $n_{i}$ and $n_{j}$ are the numbers of $i$-type and $j$-type references, respectively. This yields a continuous value between $-1$ and $1$. On one end of the spectrum, we have papers in which none of the references cites any of the other references, corresponding to a KI value of $1$ (since $n_{j}=0$). On the other extreme, we have papers in which every reference (apart from the oldest one) cites at least one other reference, corresponding to a KI value that approaches $-1$ ($\frac{1-n_{j}}{1+n_{j}}$, since the oldest reference cannot cite any of the other ones, which yields $n_{i}=1$). Fig. 1A illustrates three examples, each depicting a focal paper (represented as blue diamond) along with its references (gray circles), demonstrating how KI changes with different values of $n_{i}$ and $n_{j}$. Throughout our analyses, we exclude papers that cite only one reference, since their KI is bound to be $1$ by definition, lacking any meaningful interpretation.
@@
The quantification of KI makes it possible to systematically explore the distribution and evolution of knowledge independence in scientific papers. In general, the KI distribution seems to resemble a bell-shaped curve (Fig. 1B), punctuated by two outlier peaks at values 00 and $1$. This anomaly is due to the large number of papers with relatively short bibliographies (figure S1). For every such paper, the range of KI contains relatively few (if any) values other than 00 and $1$, e.g., for any paper with two references, $\text{KI}\in\{0,1\}$ (i.e., 00 and $1$ are the only possible values), and for any paper with four references, $\text{KI}\in\{-0.5,0,0.5,1\}$ (i.e., 00 and $1$ represent half the possible values). Additionally, $1$ is the only value that appears in the range of KI regardless of the number of references, while 00 is the only value (other than $1$) that appears in the range whenever the number of references is even. Fig. 1C depicts the evolution of KI. As can be seen, there has been a notable decline in papers’ KI over the past six decades, regardless of the discipline under consideration.
@@
In principle, building on independent knowledge creates more opportunities for novel combinations, thereby fostering innovation. To test this hypothesis, we examined the correlation between KI and the disruption index  (?, ?). The disruption index quantifies the extent to which a paper introduces novel contributions that overshadow the prior work it builds upon, leading subsequent scholars to cite the paper itself without citing its references. This index belongs to $[-1,1]$, representing a spectrum of papers ranging from fully developmental ($-1$) to fully disruptive ($1$) (see the Disruption section in materials and methods).
@@
KI is an individual measure of the extent to which a paper connects originally isolated knowledge whereas the disruption index collectively measures the peer scholars’ future appreciation of this paper. Seemingly, these two are not necessarily associated with each other. But our findings strongly support the hypothesis of innovation: papers with higher KI are more disruptive (Fig. 2A, blue). This relationship remains robust even when focusing exclusively on disruptive papers (disruption index $>0$), suggesting that increasing KI not only correlates with higher disruption scores but also raises the odds of a paper being disruptive (Fig. 2A, green). These results reveal the dual role of knowledge combinations: dependent knowledge enables incremental advancements within established bodies of research, while independent knowledge fosters innovative combinations that tend to disrupt existing scientific ideas. The pattern holds when controlling for the length of the reference list within the focal paper (figure S2), and when introducing numerous variations to the KI and disruption indices (definitions in materials and methods, and results in figure S3).
@@
Next, we control for three properties of the focal paper: impact, publication date, scientific discipline. Starting with impact, it is measured as the number of citations the paper accumulates within the first five years post-publication, denoted by $C_{5}$. As impact increases, the association between disruption and KI becomes more pronounced, especially when the KI value is small (Fig. 2B). Moving on to publication date, we find that the relationship between KI and disruption has strengthened in recent decades (Fig. 2C). We also find a systematic decline in disruption over time, consistent with prior findings  (?). Next, we consider the discipline of the focal paper. As shown in Fig. 2D, the relationship between KI and disruption holds universally, from natural to social sciences, suggesting that it is a fundamental feature of the interplay between knowledge recombination and scientific disruption.
@@
Having controlled for various properties of the focal paper, let us now control for the properties of the references within that paper. One aspect lies in the topological structure of the references, as reflected in the novelty index, capturing the degree of atypicality in the combinations of references at the journal level  (?) (details in materials and methods). The relationship between KI and disruption remains robust after controlling for novelty (figure S4B-C). Another aspect of the focal paper’s references lies in their statistical patterns, such as average age, average impact, and average disruption. In terms of reference age, both papers with high and low KI tend to cite relatively younger references compared to those cited by papers with intermediate KI values, i.e., those balancing dependent and independent knowledge (figure S5A). Furthermore, papers with higher KI exhibit a preference for less impactful (figure S6A), but more disruptive references (figure S7A). Importantly, the association between KI and disruption remains robust after controlling for reference age (figure S5B-C), reference impact (figure S6B-C), and reference disruption (figure S7B-C).
@@
We further investigate whether team composition influences the relationship between KI and disruption. We measure team size  (?), geographic distance  (?), and collaborative freshness  (?) of the focal paper (definitions in materials and methods) and then control for these variables. Our results reveal that the positive correlation between disruption and KI persists across varying team size (figure S8). Similarly, the inherent advantages of onsite and fresh teams do not alter the positive correlation between KI and disruption (figure S9-S10). These findings well dispel the concerns that team composition might confound the observed relationship, affirming the robustness of the link between KI and disruption.
@@
Motivated by the strong association between KI and disruption, we perform a matching experiment to estimate the effect size. To this end, we conduct randomized controlled trials via propensity score matching  (?) (PSM, details in materials and methods). In particular, each paper is assigned a propensity score based on a regression model accounting for covariates, allowing for the matching of papers in the treated and control groups with similar propensity scores. The effect size, characterized as the average treatment effect on the treated (ATT), is then estimated by comparing the disruption differences between matched pairs. Fig. 2E shows the pairwise ATTs across varying treated and control group sizes. The results reveal that for each controlled size, the ATT increases as the treated size grows, transiting from negative to positive when the treated size approaches the controlled size. This robustly confirms a positive effect of increasing KI on disruption.
The above findings further prompt us to ask whether KI plays a dominant role in driving disruption. To compare its influence with other variables, we employ normalized ordinary least-squares (OLS) regression (details in materials and methods) to model disruption as a function of KI alongside other variables. As summarized in Table 1, predictions from these models reveal that KI exerts a significant and dominant effect on disruption. While most other variables also show statistical significance, their impact is relatively minor. The key exception is the reference disruption, which exhibits an effect comparable to that of KI. Notably, as revealed that papers with higher KI tend to cite more disruptive references (figure S7A), suggesting that reference disruption may not be an independent driver of disruption but rather a consequence of KI. To investigate this further, we perform a mediation analysis  (?) (details in materials and methods). The results confirm that nearly half of the effect attributed to reference disruption is mediated by KI, while other variables show minimal mediating effect in this relationship compared to KI (table S1).

### v2

@@
Knowledge Independence Breeds Disruption but Limits Recognition
Xiaoyao Yu1,2 Talal Rahwan2∗ Tao Jia1,3∗
1College of Computer and Information Science
Southwest University Chongqing 400715 China.
2Computer Science
Science Division New York University Abu Dhabi Abu Dhabi 129188 UAE.
3College of Computer and Information Science
Chongqing Normal University Chongqing 401331 China.
Despite extensive research on scientific disruption, two questions remain: why disruption has declined amid growing knowledge, and why disruptive work receives fewer and delayed citations. One way to address these questions is to identify an intrinsic, paper-level property that reliably predicts disruption and explains both patterns. Here, we propose a novel measure, knowledge independence, capturing the extent to which a paper draws on references that do not cite one another. Analyzing 114 million publications, we find that knowledge independence strongly predicts disruption and mediates the disruptive advantage of small, onsite, and fresh teams. Its long-term decline, nonreproducible by null models, provides a mechanistic explanation for the parallel decline in disruption. Causal and simulation evidence further indicates that knowledge independence drives the persistent trade-off between disruption and impact. Taken together, these findings fill a critical gap in understanding scientific innovation, revealing a universal law: Knowledge independence breeds disruption but limits recognition.
Teaser
Knowledge independence, a declining intrinsic driver, mechanistically explains disruption’s paradoxes: its system-wide decline and its trade-off with impact.
Introduction
Scientific disruption has become a central topic in studies of scientific innovation  [wu2019large, park2023papers, lin2023remote, zeng2021fresh, deng2025critical, kozlov2023disruptive, leahey2023types]. Unlike most citation-based metrics  [garfield1972citation, hirsch2005index, egghe2006improvement, shen2014collective, ruiz2015field, hutchins2016relative, ke2023network, meng2024hidden], which primarily quantify the magnitude of scientific impact, the disruption index  [funk2017dynamic, wu2019large] captures the type of influence a paper exerts—specifically, whether it displaces or develops the prior work on which it builds. Despite numerous studies examining this topic, two unexplained phenomena persist. First, disruption is declining over time despite unprecedented growth in scientific output  [park2023papers] and “no one knows why”  [kozlov2023disruptive]. This puzzle is compounded by yet another unexplained phenomenon: disruption is correlated with fewer citations and delayed recognition  [bornmann2020disruption, lin2022new, liang2023bias, zeng2023disruptive, deng2025critical]. Together, these open questions constitute a fundamental gap in our understanding of scientific disruption.
One way to address these questions is to identify an intrinsic property of research papers that reliably predicts their disruption. After all, the disruptive potential of a paper must stem from something it possesses ex ante—a feature present at the moment of publication. By contrast, the disruption index, which captures how future research cites the paper, is a proxy that offers an ex post assessment: it reflects the consequences, rather than the underlying causes, of disruptive potential. Once such an intrinsic property is identified, it can shed light on the aforementioned phenomena. More specifically, if this property has declined systematically over the past decades, it would naturally account for the parallel decline in disruption. Likewise, if it turns out to be negatively associated with impact, it would serve as a confounder underlying the observed association between disruption and impact.
A promising candidate for such a property is the atypicality of the body, or “combination,” of the references on which a paper builds. This idea is grounded in recombinant growth theory, which posits as the stock of knowledge expands, so too does the combinatorial search space, creating more opportunities for unconventional pairings that can spark innovation  [weitzman1998recombinant]. One of the most established measures that capture such atypicality is the novelty index  [uzzi2013atypical], which evaluates the rarity of co-cited journal pairs in a paper’s bibliography. Intuitively, if a pair co-occurs less frequently than expected by random chance, it contributes to the paper’s atypicality (see Methods for a formal definition). There are two main approaches to representing novelty based on the rarity of journal combinations: one based on the median rarity of distribution  [lin2022new, park2023papers], and another based on the left tail (10th percentile) of that distribution  [uzzi2013atypical, lin2023sciscinet]. Notably, these two variants lead to contradictory conclusions: the median-based approach yields a positive association between novelty and disruption  [lin2022new], whereas the tail-based alternative produces a negative association  [li2025innovation]. This divergence indicates that the association is highly sensitive to the operationalization of novelty, and hence this index cannot serve as the intrinsic property we seek.
An alternative way to assess a paper’s atypicality focuses on the disciplines, not journals, that the paper combines in its bibliography  [rafols2010diversity, zhang2016diversity, zhang2021scientometric, xiang2025evaluating]. Arguably, the most establish measure in this line of work is the Rao–Stirling interdisciplinarity index  [stirling2007general, rafols2010diversity, zhang2016diversity, zhang2021scientometric, xiang2025evaluating]. Yet the empirical evidence linking interdisciplinarity to disruption is inconsistent. An analysis based on the Scopus dataset reported a positive association between interdisciplinarity and disruption  [chen2024interdisciplinarity]; a study using the Microsoft Academic Graph dataset found the association to be negative  [liu2024monodisciplinary]; and a third study using the Web of Science dataset observed a negative association with the disruption index but a positive association with the odds of a paper being disruptive  [cheng2025impact]. Taken together, these mixed findings indicate that the relationship between interdisciplinarity and disruption is highly sensitive to the dataset and to the particular operationalization of disruption, suggesting that interdisciplinarity, like novelty, cannot serve as the robust predictor we are looking for.
Since journal- and discipline-based atypicality indices cannot reliably predict disruption, a natural question follows: What specific form of atypical reference combinations, if any, can do so—and might such a form also clarify the broader phenomena surrounding disruption? To address this question, we turn to a largely overlooked aspect of knowledge combination: path dependence. Here, by path dependence we do not refer to the traditional sociological concept of historical lock-in and self-reinforcement  [page2006path, strambach2013reconceptualizing, puffert2024path], but rather to the extent to which technological innovations rely on established sequences, or “paths,” of prior inventions  [xie2016does, zhang2023knowledge]. From the latter perspective, studies have shown that reliance on established technological trajectories tends to reinforce incremental improvements by reducing search and development costs  [ahuja2001entrepreneurship, katila2002something, kapoor2012firms], and hinders radical innovation by narrowing the exploratory space  [kaplan2015double, xie2016does, schillebeeckx2021knowledge]. What remains unknown, however, is whether similar dynamics extend from technological innovation to scientific disruption. Unfortunately, existing measures of this notion of knowledge path dependence are either qualitative  [sydow2009organizational, jackson2010formulation, jackson2012modeling] or defined at the institutional- or regional-level  [jaffe1986technological, zhang2023knowledge], leaving us without a quantitative, paper-level metric capable of capturing path dependence in scientific research
Here, we propose a measure that quantifies the extent to which a paper’s references are “independent.” More concretely, it calculates the proportion of references that are independent from, i.e., do not cite, other references in the paper’s bibliography. Consider, for instance, a pair of references $(r_{1},r_{2})$ in the focal paper’s bibliography. If it turns out that $r_{1}$ cites $r_{2}$, then this can be thought of as a path transmitting knowledge from $r_{2}$ to $r_{1}$. As such path-dependent references become more frequent, knowledge path independence decreases. For brevity, we hereafter omit the term “path” and simply write “knowledge independence,” denoted by KI. Note that our measure captures the independence, rather than dependence, of knowledge paths, since we are primarily interested in atypical, rather than typical, combinations. Conceptually, a paper built on independent knowledge acts as a “knowledge broker” connecting intellectual entities that were previously unlinked. This is quite different from the role played by papers that are atypical from the perspective of novelty (combining unconventional journal pairs) or interdisciplinarity (combining divergent disciplines); those papers act as “knowledge importers” bridging distant knowledge domains. Importantly, KI does not necessarily correlate with novelty or interdisciplinarity. For instance, a paper can have low novelty but high KI, e.g., if it cites papers from two commonly co-cited journals, but its references hardly cite each other. Likewise, a paper may have high interdisciplinarity but low KI, e.g., if it links divergent disciplines, but many of its references happen to cite one another.
Armed with our measure, we examine 53.8 million publications from the Web of Science, and find that KI positively correlates with disruption. Causal inference using standard matching techniques (CEM and PSM) along with normalized regression analyses reveal that this association is significant. Moreover, both the effect magnitude and the explanatory power of KI on disruption are substantially greater than that of numerous intrinsic properties, including the paper’s discipline, the publication year, the team composition, and various reference attributes. Our analysis also reveals KI’s mediating role in the well-documented association between team composition and disruption. More specifically, we show that small  [wu2019large], onsite  [lin2023remote], and fresh  [zeng2021fresh] teams exhibit a systematic preference for higher KI, which helps explain their disruptive potential. Taken together, these findings highlight KI as the intrinsic paper property that most strongly predicts disruption. Finally, we show that KI provides a plausible unified explanation for the aforementioned paradoxes surrounding disruption. In particular, we find that KI has undergone a decades-long decline across all disciplines—a trend that cannot be replicated by null models—signaling a genuine shift in scientists’ behaviour, and explaining the parallel decline in disruption. Furthermore, we observe that papers built on higher KI experience lower and delayed citation recognition. Once KI, the confounder, is controlled for, the negative correlation between disruption and citation recognition is substantially attenuated, providing a plausible explanation for the second paradox. More broadly, our findings fill a critical gap in our understanding of scientific innovation, revealing a universal law: Knowledge independence breeds disruption but limits recognition.
Results
@@
To quantify knowledge independence in the context of bibliometrics, we develop a measure that relies on the citation network. This is because citations serve as tangible indicators of knowledge flow, reflecting the dependency of the citing paper on its references. Intuitively, for any given paper, the measure quantifies the degree to which its references are “independent” in the sense that they do not cite one another. To this end, let us introduce two types of references: An indind-type reference is one that does not cite any other work within the same reference list, whereas a depdep-type reference is one that does cite at least one other work in this list. Then, knowledge independence (KI) is measured as the difference between the fraction of indind-type and depdep-type references. More formally:
$\text{KI}=\frac{n_{\rm ind}-n_{\rm dep}}{n_{\rm ind}+n_{\rm dep}},$ (1)
@@
where $n_{\rm ind}$ and $n_{\rm dep}$ are the numbers of indind-type and depdep-type references, respectively. This yields a continuous value between 1-1 and $1$. On one end of the spectrum, we have papers in which none of the references cites any of the other references, corresponding to a KI value of $1$ (since $n_{\rm dep}=0$). On the other extreme, we have papers in which every reference (apart from the oldest one) cites at least one other reference, corresponding to a KI value that approaches 1-1 ($\frac{1-n_{\rm dep}}{1+n_{\rm dep}}$, since the oldest reference cannot cite any of the other ones, which yields $n_{\rm ind}=1$). Fig. 1A illustrates three examples, each depicting a focal paper (represented as blue diamond) along with its references (gray circles), demonstrating how KI changes with different values of $n_{\rm ind}$ and $n_{\rm dep}$. Throughout our analyses, we exclude papers that cite only one reference, since their KI is bound to be $1$ by definition, lacking any meaningful interpretation.
Figure 1: $|$ Quantifying knowledge independence. (A) Three examples of how knowledge independence (KI) is calculated for a given focal paper (blue diamond) based on its references (gray circles). A reference is of indind-type (green) if it does not cite any other reference, or of depdep-type (orange) if it does. KI is then the difference in the fraction of indind-type and depdep-type references. (B) The KI distribution of 73,256,108 papers with at least two references published between 1950 and 2024 in the OpenAlex. (C) The pervasive downtrend of KI over time across disciplines. Bootstrapped 95% confidence intervals are shown as shaded bands.
@@
The quantification of KI makes it possible to systematically explore the distribution and evolution of knowledge independence in scientific papers. In general, the KI distribution seems to resemble a bell-shaped curve (Fig. 1B), punctuated by two outlier peaks at values 0 and $1$. This anomaly is due to the large number of papers with relatively short bibliographies (Extended Data Fig. 1). For every such paper, the range of KI contains relatively few (if any) values other than 0 and $1$, e.g., for any paper with two references, $\text{KI}\in\{0,1\}$ (i.e., 0 and $1$ are the only possible values), and for any paper with four references, $\text{KI}\in\{-0.5,0,0.5,1\}$ (i.e., 0 and $1$ represent half the possible values). Additionally, $1$ is the only value that appears in the range of KI regardless of the number of references, while 0 is the only value (other than $1$) that appears in the range whenever the number of references is even. Fig. 1C depicts the evolution of KI. As can be seen, there has been a notable decline in papers’ KI over the past six decades, regardless of the discipline under consideration.
To validate the KI metric and its observed patterns in distribution and temporal trend, we conducted a two-stage robustness analysis. First, we show through direct statistical controls that our findings are robust to variations in reference count, novelty index, and interdisciplinarity index (see Section S1 in Supplementary Materials). More fundamentally, to ensure these patterns are not artifacts of network evolution, we employed a random rewiring Monte Carlo simulation. The results reveal that the empirically observed KI has a distribution and historical decline that are starkly different from the random expectation (see Section S2.1 in Supplementary Materials). This provides strong evidence that our findings reflect a substantive, non-random shift in scientific behavior.
@@
In principle, building on independent knowledge creates more opportunities for unconventional combinations, thereby fostering innovation with a greater chance. To test this hypothesis, we examined the correlation between KI and the disruption index  [funk2017dynamic, wu2019large]. The disruption index quantifies the extent to which a paper introduces novel contributions that overshadow the prior work it builds upon, leading subsequent scholars to cite the paper itself without citing its references. This index belongs to [1,1][-1,1], representing a spectrum of papers ranging from fully developmental (1-1) to fully disruptive ($1$) (see the Disruption section in Methods).
@@
KI is an individual measure of the extent to which a paper connects originally isolated knowledge within the internal reference network, whereas the disruption index is a collective measure of future peer recognition for a paper’s contribution relative to its predecessors. Put it differently, the former measure is solely rooted in the cited bibliography at the publication moment, but the latter index integrates the past context (cited bibliography) with the future recognition patterns (received citations). Seemingly, a paper’s intrinsic structural property (KI) and its future extrinsic recognition (disruption) are not necessarily associated with each other. But our findings strongly support the hypothesis of innovation: papers with higher KI are more disruptive (Fig. 2A, blue). This relationship remains robust even when focusing exclusively on disruptive papers (disruption index >0>0), suggesting that increasing KI not only correlates with higher disruption scores but also raises the odds of a paper being disruptive (Fig. 2A, green). This pattern is not an artifact of reference count, as it holds when controlling for the reference count within the focal paper (Extended Data Fig. 2); Our observations are also robust to the variations of KI and disruption indices (definitions in Methods, and results in Extended Data Fig. 3).
Figure 2: $|$ Knowledge independence is associated with scientific disruption. (A) For the 55,576,192 papers published before 2023 in the OpenAlex that are cited at least once, the average disruption percentile (blue curve, left y-axis) and the disruption positive ratio (green curve, right y-axis) increase with the KI percentile. (B) The association between KI and disruption persists, regardless of the impact of the focal paper. Here, impact is measured as the number of citations received within the first five years after publication, denoted by $C_{5}$. (C) The association between KI and disruption persists across decades. (D) The association between KI and disruption persists, regardless of discipline. Bootstrapped 95% confidence intervals are shown as shaded bands in panels A-D. (E) The ATT (average treatment effect on the treated) matrix of KI on disruption via coarsened exact matching (CEM). Each controlled group is set as a baseline, and ATTs are calculated for comparisons between the baseline and each of the treated groups. Blue cells represent negative ATTs, while red ones represent positive ATTs, with color intensity proportional to the absolute value. Within each controlled group, the ATT increases as the treated size grows, transitioning from negative to positive when the treated size approaches the controlled size. Each ATT is tested against the null hypothesis (ATT equals 0) using a two-sided t-test. (*$p<0.05$, **$p<0.01$, ***$p<0.001$).
@@
Next, we control for three properties of the focal paper: impact, publication date, scientific discipline. Starting with impact, it is measured as the number of citations the paper accumulates within the first five years post-publication, denoted by $C_{5}$. As impact increases, the association between disruption and KI becomes more pronounced, especially when the KI value is small (Fig. 2B). Moving on to publication date, we find that the relationship between KI and disruption has strengthened in recent decades (Fig. 2C). We also find a systematic decline in disruption over time, consistent with prior findings  [park2023papers]. Next, we consider the discipline of the focal paper. As shown in Fig. 2D, the relationship between KI and disruption holds universally, from natural to social sciences, suggesting that it is a fundamental feature of the interplay between knowledge recombination and scientific disruption.
@@
Having controlled for various properties of the focal paper, let us now control for the properties of the references within that paper. One aspect lies in atypical combinations of references at the journal and disciplinary level, as reflected in the novelty index and the interdisciplinarity index, respectively (details in Methods). The relationship between KI and disruption remains robust after controlling for both indices (Extended Data Fig. 4B-C, Extended Data Fig. 5B-C). This addresses a critical concern: while established atypicality metrics (novelty and interdisciplinarity) primarily measure the conceptual distance based on meso-level co-occurrence (journals or disciplines), KI captures a distinct and more fine-grained micro-level structural facet: the direct independence of the relationship among papers. KI’s robust and independent predictive power—even after controlling for these rivals—confirms that this structural dimension is a unique and non-redundant component. Another aspect of the focal paper’s references lies in their statistical patterns, such as average age and impact. In terms of reference age, both papers with high and low KI tend to cite relatively younger references compared to those cited by papers with intermediate KI values, i.e., those balancing dependent and independent knowledge (Extended Data Fig. 6A). As for reference impact, papers with higher KI exhibit a preference for less impactful (Extended Data Fig. 7A). Importantly, the association between KI and disruption remains robust after controlling for reference age (Extended Data Fig. 6B-C) and reference impact (Extended Data Fig. 7B-C).
@@
We further investigate whether team composition influences the relationship between KI and disruption. We measure team size  [wu2019large], geographic distance  [lin2023remote], and collaborative freshness  [zeng2021fresh] of the focal paper (definitions in Methods) and then control for these variables. Our results reveal that the positive correlation between disruption and KI persists across varying team size (Extended Data Fig. 9). Similarly, the inherent advantages of onsite and fresh teams do not alter the association between KI and disruption (Extended Data Fig. 10-11). These findings well dispel the concerns that team composition might confound the observed relationship, affirming the robustness of the link between KI and disruption.
@@
Motivated by the strong association between KI and disruption, we perform a matching experiment to estimate the effect size. To this end, we conduct randomized controlled trials via Coarsened Exact Matching  [blackwell2009cem, iacus2012causal] (CEM, details in Methods). In particular, each ex-ante covariate is coarsened into bins with appropriate scheme, allowing for the exact matching within these coarsened strata; comprehensive robustness analyses on different coarsening schemes are displayed in Supplementary Materials (Section S3). The effect size, characterized as the average treatment effect on the treated (ATT), is then estimated by a weighted regression model on the matched sample using the CEM-generated weights. Fig. 2E shows the pairwise ATTs across varying treated and control group sizes. The results reveal that for each controlled size, the ATT increases as the treated size grows, transiting from negative to positive when the treated size approaches the controlled size. This robustly confirms a positive effect of increasing KI on disruption. To ensure our causal estimates are robust to the matching algorithm, we supplemented our primary CEM analysis with Propensity Score Matching (PSM). This additional analysis yields fully consistent results, reinforcing the significant positive impact of KI on disruption (see Section S4 in Supplementary Materials).
Knowledge independence dominantly predicts disruption and explains its decline
The above analyses establish KI as a robust predictor of disruption. What remains unknown is whether KI can resolve the explanatory gap of atypicality metrics? To find out, we employed normalized ordinary least-squares (OLS) regression to competitively test KI against all other covariates.
Table 1: $|$ Normalized regression models of disruption and the influencing factors.
Model 1 Model 2 Model 3 Model 4 Model 5 Model 6 Model 7 Model 8 Model 9 $R^{2}$%
KI 0.4389*** 0.4389*** 0.4409*** 0.4178*** 0.4399*** 0.4382*** 0.4364*** 0.4335*** 0.4134*** 71.87
(0.000) (0.000) (0.000) (0.000) (0.000) (0.000) (0.000) (0.000) (0.000)
Novelty 0.0001 0.0002 0.00
(0.000) (0.000)
Interdisciplinarity -0.0120*** -0.0061*** 0.61
(0.000) (0.000)
Reference count -0.0623*** -0.0579*** 10.82
(0.000) (0.000)
Reference age -0.0148*** -0.0148*** 0.12
(0.000) (0.000)
Team size -0.0147*** -0.0159*** 0.68
(0.000) (0.000)
Team distance -0.0307*** -0.0256*** 1.21
(0.000) (0.000)
Team freshness 0.0470*** 0.0469*** 2.45
(0.000) (0.000)
Discipline fixed effects Yes Yes Yes Yes Yes Yes Yes Yes Yes 10.60
Year fixed effects Yes Yes Yes Yes Yes Yes Yes Yes Yes 1.64
N 33304580 33304580 33304580 33304580 33304580 33304580 33304580 33304580 33304580
$R^{2}$ 0.225 0.225 0.225 0.228 0.225 0.225 0.226 0.227 0.231
  • $\dagger$
    We perform normalized ordinary-least-squares (OLS) regression analyses to examine the effect of KI on disruption, while controlling for various covariates. Each regression coefficient is tested against the null hypothesis (coefficient equals 0) using a two-sided t-test. Standard errors are provided in parentheses for each coefficient. Note that we do not apply adjustments for multiple hypothesis testing in this analysis. (*$p<0.05$, **$p<0.01$, ***$p<0.001$).
The results are conclusive. In the full model (Table 1, Model 9) accounting for all ex-ante variables, KI maintains its large, positive, and significant association with disruption ($0.4134$), with a dominant share in accounting for $R_{2}$ ($71.87\%$). In contrast, the atypicality metrics, both novelty index and interdisciplinarity display a minor association ($0.0002$ and 0.0061-0.0061) with low share in $R_{2}$ ($0.00\%$ and $0.61\%$). This finding persists when applied to alternative disruption metrics (Table S1 in Section S5), suggesting that the key to disruptive innovation is not the combination of conceptually distant knowledge (novelty or interdisciplinarity), but the integration of independent knowledge (KI).
KI’s dominant predicting power provides an immediate and direct mechanistic explanation for the first paradox, i.e., the systematic decline in disruption  [park2023papers]. As shown in Fig. 1C, KI itself has undergone a steady, decades-long decline across all disciplines. Meanwhile, we find that the relationship between KI and disruption has strengthened in recent decades and accompanied with a systematic decline in disruption over time (Fig. 2C), consistent with prior findings  [park2023papers]. One might argue that this parallel decline could be caused by mere chance or by the structural evolution of the citation network itself. We rule out this possibility with our randomized rewiring Monte Carlo analysis (see Section S2.1 in Supplementary Materials): the randomized network fails to reproduce the decline trend of KI, signaling that it is a consequence of a behavioural shift in the way knowledge is combined. The randomized network also fails to replicate the KI-disruption association, further indicating that the phenomenon we observe is derived from non-trivial scientists’ behaviour. This temporal alignment—where the intrinsic property (KI) and the extrinsic outcome (disruption) decline in parallel—provides a plausible explanation for why disruption is decreasing.
Finally, we extended this analyzing framework to the technological domain. Results from the OECD patent dataset consistently show that a higher KI is also a key driver of technological disruption (see Section S8 in Supplementary Materials). The consistency of these results across the scientific and technological domains strongly reinforces the predictive power of the KI metric on disruption.